# DQ-DETR: DETR with Dynamic Query for Tiny Object Detection

### v1

@@ Abstract
Despite previous DETR-like methods having performed successfully in generic object detection, tiny object detection is still a challenging task for them since the positional information of object queries is not customized for detecting tiny objects, whose scale is extraordinarily smaller than general objects. Also, DETR-like methods using a fixed number of queries make them unsuitable for aerial datasets, which only contain tiny objects, and the numbers of instances are imbalanced between different images. Thus, we present a simple yet effective model, named DQ-DETR, which consists of three different components: categorical counting module, counting-guided feature enhancement, and dynamic query selection to solve the above-mentioned problems. DQ-DETR uses the prediction and density maps from the categorical counting module to dynamically adjust the number of object queries and improve the positional information of queries. Our model DQ-DETR outperforms previous CNN-based and DETR-like methods, achieving state-of-the-art mAP 30.2% on the AI-TOD-V2 dataset, which mostly consists of tiny objects.
@@ 1 Introduction
Nevertheless, in the previous DETR-like methods, the object queries used in the transformer decoder do not consider the number and position of instances in the image. Generally, they apply the fixed number K of object queries, where K represents the maximum number of the detection objects, e.g., K=100, 900 in DETR and DINO-DETR, respectively. In addition, the position of object queries is a set of learned embeddings, which are irrelevant to the current image and do not have explicit physical meaning to tell where the queries are focusing on.
We argue that the previous DETR-like methods are inappropriate in aerial images, which only contain tiny objects and have an imbalance of object instances between different images. For example, in the AI-TOD-V2 dataset [24], some images have more than 1500 objects, but others have less than 10 objects. We observe that using smaller K restricts the recall of the objects in dense images, leaving lots of instances undetected (FN), as shown in Fig. 1(a). Conversely, Fig. 1(c) shows using larger K in the sparse images not only introduces many underlying false positive samples (FP) but also causes a waste of computing resources since the computing complexity in the decoder’s self-attention modules grows quadratic with the number of queries K.
Stemming on the above-mentioned weakness, we proposed a novel DETR-like method named DQ-DETR. In this work, we propose a dynamic query selection module for adaptively choosing different numbers of object queries in DETR’s decoder stage, resulting in fewer FP in sparse images and fewer FN in dense images. Moreover, we generate the density maps and estimate the number of instances in an image by the categorical counting module. The number of object queries is adjusted based on the predicted counting number. In addition, we aggregate the density maps with the visual feature from the transformer encoder to reinforce the foreground features, enhancing the spatial information for tiny objects. The strengthened visual feature will be further used to improve the positional information of object queries. As such, we can simultaneously handle the images with few and crowded tiny objects by dynamically adjusting the number of object queries used in the decoder.
@@ 2 Related work
DETR-like Methods
[2] proposes an end-to-end object detection framework based on the transformer named DETR (DEtection TRansformer), where the transformer encoder extracts instance-level features from an image, and the transformer decoder uses a set of learnable queries to probe and pool features from images. While DETR achieves comparable results with the previous classical CNN-based detectors [16, 19], it suffers severely from the problem of slow training convergence, needing 500 epochs of training to perform well. Many follow-up works have attempted to address the slow training convergence of DETR from different perspectives.
@@
Tiny Object Detection.
@@
However, these methods highly rely on the predefined threshold, which is unstable to a different dataset. Instead, our DQ-DETR uses a counting-guided feature enhancement module to improve the visual features with more spatial information about tiny objects’ scale and location. Also, although the above-mentioned DETR-like methods have improved the formulation of queries, they are not specially designed to detect tiny objects. For example, the object queries in [10, 12, 2, 7] are learned from the training data and remain the same for different input images. We argue that the static positions of object queries are unsuitable for aerial datasets, where the distribution of instances varies extremely in different images, i.e., some images contain dense objects concentrated in specific areas, while some only have a few objects scattered throughout the images. Our proposed DQ-DETR is the first DETR-like model that focuses on tiny object detection. DQ-DETR dynamically adjusts the number of object queries and enhances the position information of queries in order to detect tiny objects precisely.
@@ 3.3 Categorical Counting Module
Counting Number Classification.
Lastly, we estimate the counting number $N$, i.e., the number of instances per image, by a classification head and categorize them into four levels, which are $N\leq 10$, $10<N\leq 100$, $100<N\leq 500$, and $N>500$. The classification head consists of two linear layers. Further, the numbers 10, 100, and 500 are selected based on the dataset’s characteristics, i.e., the mean and standard deviation of the number of instances $N$ per image. Notably, we do not use the regression head as in the traditional crowd-counting methods, which regresses the counting number to a specific number. We attribute the reason to the drastic difference of the number of instances in each image, where $N$ ranges from 1 to 2267 in different images of AI-TOD-V2. It is difficult to regress an accurate number, hurting the detection performance (refer to Sec 4.4.3 Table 5).
@@ 3.4 Counting-Guided Feature Enhancement Module (CGFE)
We refine EMSV feature with the density map from categorical counting module through the proposed Counting-Guided Feature Enhancement Module (CGFE) to improve the spatial information of tiny objects. In addition, the refined features will further used to enhance the position information of queries. This module comprises a 2-D spatial attention module and a 1-D channel attention module.
Spatial attention map.
We employ a $1\times 1$ convolution layers to down-sample the density map ${F}_{c}$, creating multi-scale counting feature maps ${F}_{c,i}\in\{1,2,\dots,l\}$ to in line with the shape of each layer of encoder’s multi-scale feature maps $S_{i}\in\{1,2,\dots,l\}$. Subsequently, we first apply average pooling (AvgP.) and max pooling (MaxP.) on each layer of multi-scale counting features ${F}_{c,i}\in\mathbb{R}^{b\times 256\times h\times w}$ along the channel axis. Then, the two pooling features $\mathbb{R}^{b\times 1\times h\times w}$ are concatenated and sent into a 7x7 convolution layer followed by a Sigmoid function to produce spatial attention map ${W}_{s}\in\mathbb{R}^{b\times 1\times h\times w}$. We formulate this process in Eq. 1.
@@
Channel attention map.
After the spatial attention, we further apply channel attention to the spatial-intensified features ${E}_{i}$, exploiting the inter-channel relationship of features. Specifically, we first apply average pooling and max pooling on each layer of ${E}_{i}\in\mathbb{R}^{b\times 256\times h\times w}$ along the spatial dimension. Next, the two pooling features $\mathbb{R}^{b\times 256\times 1\times 1}$ are sent into a shared MLP and merged together with element-wise addition to create channel attention map ${W}_{c,i}$. Finally, the channel attention map ${W}_{c,i}\in\mathbb{R}^{b\times 256\times 1\times 1}$ is multiplied with original $E_{i}\in\mathbb{R}^{b\times 256\times h\times w}$ and further get the counting-guided intensified feature maps ${F}_{t}$. The formulas are defined in Eq. 3 and Eq. 4:
${W}_{c,i}=\sigma(MLP(AvgP.({E}_{i}))+$ (3)
$MLP(MaxP.({E}_{i}))),$
@@ 4 Experiments
{tblr}
columneven = c, column3 = c, column5 = c, column7 = c, column9 = c, cell21 = c=10c, cell111 = c=10c, hline1-3,11-12,17 = -, Method & #Epochs Backbone $\mathrm{AP}$ $\mathrm{AP_{50}}$ $\mathrm{AP_{75}}$ $\mathrm{AP_{vt}}$ $\mathrm{AP_{t}}$ $\mathrm{AP_{s}}$ $\mathrm{AP_{m}}$
@@
{tblr}
cells = c, hline1-2,5 = -, vline2-4 = -, Categorical Counting Dynamic Query Selection Feature Enhancement $\mathrm{AP}$ $\mathrm{AP_{vt}}$ $\mathrm{AP_{t}}$ $\mathrm{AP_{s}}$ $\mathrm{AP_{m}}$
@@
✓ ✓ 28.1 12.3 27.8 34.6 44.1
✓ ✓ ✓ 30.2 15.3 30.5 36.5 44.6
4.1 Dataset
AI-TOD-V2 This dataset includes 28,036 aerial images with 752,745 annotated object instances. There are 11,214 images for the train set, 2,804 for the validation set, and 14,018 for the test set. The average object size in AI-TOD-V2 is only 12.7 pixels, with 86% of objects in the dataset smaller than 16 pixels, and even the largest object is no bigger than 64 pixels. Also, the number of objects in an image can vary enormously from 1 to 2667, where the average number of objects per image is 24.64 with a standard deviation of 63.94.
Evaluation Metric We use the AP (Average Precision) metric with a max detection number of 1500 to evaluate the performance of our proposed method. Specifically, AP means the average value from $\mathrm{AP_{0.5}}$ to $\mathrm{AP_{0.95}}$ , with IoU interval of 0.05. Moreover, $\mathrm{AP_{vt}}$, $\mathrm{AP_{t}}$, $\mathrm{AP_{s}}$, and $\mathrm{AP_{m}}$ are for very tiny, tiny, small, and medium scale evaluation in AI-TOD [21].
4.2 Implementation Details
Based on DETR-like structure, we use a 6-layer transformer encoder, a 6-layer transformer decoder with 256 as the hidden dimension, and a ResNet50 as our CNN backbone. Furthermore, we train our model for 24 epochs with Adam optimizer and weight decay 0.0001 using 2 NVIDIA 3090 GPUs. The batch size is set to 1 due to memory constraints. For the learning rate scheduler, the initial learning rate (lr) is 0.0001, and lr is dropped at 13-th and 21-st epoch by multiplying 0.1. We use the same random crop and scale augmentation strategies following DETR [2]. Also, we apply a two-stage training scheme. We first train the categorical counting module to achieve more stable results for the number of queries in the transformer decoder. After the counting result is stabilized, we add the counting-guided feature enhancement module into training to refine the encoder’s visual features with density maps.
4.3 Main Results
Table 4 shows our main results on the AI-TOD-V2 test split. We compare the performances of our DQ-DETR with strong baselines, including both CNN-based and DETR-like methods. All CNN-based methods except YOLOv3 use ResNet50 as the backbone with feature pyramid network (FPN) [8]. Moreover, since there is no previous research of DETR-like models on tiny object detection, our DQ-DETR is the first DETR-like model that focus on detecting tiny objects. We re-implementing a series of DETR-like models on AI-TOD-V2, and all DETR-like methods except DETR use 5 scales of feature maps with deformable attention [28]. For 5-scale feature maps, features are extracted from stages 1, 2, 3, and 4 of the backbone, and add the extra feature by down-sampling the output of stage 4.
The results are summarized in Table 4, our proposed DQ-DETR achieves the best result 30.2 AP compared with other state-of-the-art methods, including CNN-based and DETR-like methods. Also, DQ-DETR surpasses the baseline by 20.5%, 20.6%, 14.1%, and 12.3% in terms of $\mathrm{AP_{vt}}$, $\mathrm{AP_{t}}$, $\mathrm{AP_{s}}$, $\mathrm{AP_{m}}$. The performance gain is greater on $\mathrm{AP_{vt}}$, and $\mathrm{AP_{t}}$, and our DQ-DETR outperforms the advanced series of DETR-like models on AI-TOD-V2. We credit the performance gains for the following reasons: (1) DQ-DETR fuses the transformer visual features with a density map from the categorical counting module to improve the positional information of object queries, which makes the queries more suitable for localizing tiny objects. (2) Our dynamic query selection adaptively chooses the number of object queries used for the detection task and can handle the images with either few or crowded objects.
{tblr}
column3 = c, column4 = c, column5 = c, column6 = c, column7 = c, column8 = c, column9 = c, column10 = c, vline2-4 = -, hline1-2,7,12 = -, Model #objects in image #Query $\mathrm{AP}$ $\mathrm{AP_{50}}$ $\mathrm{AP_{75}}$ $\mathrm{AP_{vt}}$ $\mathrm{AP_{t}}$ $\mathrm{AP_{s}}$ $\mathrm{AP_{m}}$
@@ 4.3 Main Results
{tblr}
columneven = c, column3 = c, column5 = c, vline2-5 = -, hline1-2,6-7 = -, #Objects in image Accuracy(%) AP (DQ-DETR) AP (Baseline) #Sample
@@
4.4 Ablation Study
Categorical counting module, counting-guided feature enhancement, and dynamic query selection are the newly proposed contributions in this paper. We conduct a series of ablation studies with ResNet50 backbone trained for 24 epochs on AI-TODv2 trainval set and test on AI-TODv2 test set to verify the effectiveness of each component proposed in this paper. DINO-DETR is chosen as the comparing DETR-like baseline.
4.4.1 Main ablation experiment
Table 2 shows the performance of our contributions separately on AI-TODv2. The results demonstrate that each component in DQ-DETR contributes to performance improvement. We attain an improved +2.2 AP over the baseline with the counting module and dynamic query selection. Furthermore, our counting-guided feature enhancement module fuses the encoder’s feature with spatial information from counting features. It gains an extra improvement of +4.3, +2.6, +5.2 on $\mathrm{AP}$, $\mathrm{AP_{vt}}$, and $\mathrm{AP_{t}}$ over the baseline. Consequently, we prove the power of each component in DQ-DETR on AI-TOD-V2.
4.4.2 Ablation of DQ-DETR with different number of instances in images
We explore our DQ-DETR’s performance with different numbers of instances in an image. We classify the AI-TOD-V2 dataset into 4 levels based on the number of instances $N$ in the image as in the categorical counting module, i.e., $N\leq 10$, $10<N\leq 100$, $100<N\leq 500$, and $500<N$. Our proposed DQ-DETR’s performance is analyzed under these four situations. The results are shown in Table 4.3 compared with DINO-DETR as the baseline. Our DQ-DETR dynamically adjusts the number of object queries based on the number of instances in the image. while DINO-DETR always uses 900 queries in all situations.
We can observe that in the situations of $N\leq 10$, and $10<N\leq 100$, our DQ-DETR uses fewer numbers of queries and outperforms the baseline by 16%, and 16.4% in terms of $\mathrm{AP}$. The performances in terms of $\mathrm{AP_{vt}}$, $\mathrm{AP_{t}}$ surpass the baseline by 19.8%, and 20.8% as well. Moreover, it is noteworthy that DINO-DETR performs poorly when $N>500$. Under these circumstances, there might be over 900 instances in some of the images, which is beyond the detection capability of DINO-DETR. In dense images, the detection limitation of DINO-DETR with only 900 queries will lead to many objects undetected (FN), causing a lower $\mathrm{AP}$. Our DQ-DETR dynamically selects more queries for dense images, remarkably surpassing the baseline by 42.1% in terms of $\mathrm{AP_{vt}}$.
4.4.3 Ablation of Categorical Counting Module
{tblr}
column2 = c, column3 = c, column4 = c, column5 = c, column6 = c, vline2 = -, hline1-2,5 = -, Method $\mathrm{AP}$ $\mathrm{AP_{vt}}$ $\mathrm{AP_{t}}$ $\mathrm{AP_{s}}$ $\mathrm{AP_{m}}$
@@ 4.4.3 Ablation of Categorical Counting Module
Table 5 compares our DQ-DETR’s performance of using classification or regression in the categorical counting module. The traditional crowd-counting methods usually regress the predicted counting number to a specific value. However, in our study, we use a classification head instead. This experiment demonstrates our DQ-DETR performance with these two methods. For the classification task, we classify the images into 4 classes and apply different numbers of queries in the transformer decoder, as we mentioned in the previous section. For the regression task, we regress an integer directly to predict the number of objects in the images.
The results demonstrate that using regression as a counting method performs extremely poorly. We impute the drastic performance drop for the following reasons: (1) It is challenging to regress an accurate number since the number of instances per image may vary significantly from 1 to 2267 in the AI-TOD-V2 dataset. (2) The unstable regression result will largely affect the number of queries used in the transformer decoder and cause the DETR model hard to converge.
Owing to the above reasons, we believe that classifying how many objects exist in the image into different levels is a simpler way in contrast to regression. Thus, we take classification instead of regression as a preferable method in our proposed categorical counting module.

### v2

@@ Abstract
Despite previous DETR-like methods having performed successfully in generic object detection, tiny object detection is still a challenging task for them since the positional information of object queries is not customized for detecting tiny objects, whose scale is extraordinarily smaller than general objects. Also, DETR-like methods using a fixed number of queries make them unsuitable for aerial datasets, which mostly contain tiny objects, and the numbers of instances are imbalanced between different images. Thus, we present a simple yet effective model, DQ-DETR, consisting of three components: categorical counting module, counting-guided feature enhancement, and dynamic query selection to solve the above-mentioned problems. DQ-DETR uses the prediction and density maps from the categorical counting module to dynamically adjust the number and positional information of object queries. Our model DQ-DETR outperforms previous CNN-based and DETR-like methods, achieving state-of-the-art mAP 30.2% on the AI-TOD-V2 dataset, which mostly consists of tiny objects.
Keywords DETR  $\cdot$ Tiny Object Detection
@@ 1 Introduction
Table 1: Comparison of DETR-like models’ queries strategy under different situations.
Sparse Dense Imbalance Characteristics
Deformable DETR [4] Sparse Queries (K=300) with One-To-One Assignment; Low Recall
DDQ-DETR [8] Dense Distinct Queries (K=900); Low Recall if #Object $\gg$ #Query
DQ-DETR(Ours) Dynamically adjust the "Number" and "Position" of Queries
We argue that the previous DETR-like methods are inappropriate in aerial images, which only contain tiny objects and have an imbalance of instances between different images. In the previous DETR-like methods, the object queries used in the transformer decoder do not consider the number and position of instances in the image. Generally, they apply the fixed number K of object queries, where K represents the maximum number of the detection objects, e.g., K=100, 900 in DETR and DINO-DETR, respectively. DETR [3] and Deformable-DETR [4] apply a fixed number of sparse queries, suffering a low recall rate. To address the problem, DDQ [8] select dense distinct queries, K=900, with a class-agnostic NMS based on a hand-designed IoU threshold. Though DDQ applies dense queries for detection, the number of queries is still limited.
The fixed number of queries will cause low detection accuracy in aerial datasets since the number of objects may vary drastically in different images. In the AI-TOD-V2 dataset [9], some images have more than 1500 objects, but others have less than 10 objects. Under the situation that the number of objects in images is more than the DETR’s query number K, a low recall rate is an expected issue. Using smaller K restricts the recall of the objects in dense images, leaving lots of instances undetected (FN). While using a large K in the sparse images not only introduces many underlying false positive samples (FP) but also causes a waste of computing resources since the computing complexity in the decoder’s self-attention modules grows quadratic with the number of queries K.
Besides, in the previous DETR-like methods, the object queries used in the transformer decoder do not consider the position of instances in the image. The position of object queries is a set of learned embeddings, which are irrelevant to the current image and do not have explicit physical meaning to tell where the queries are focusing on. The static positions of object queries are unsuitable for aerial datasets, where the distribution of instances varies extremely in different images, i.e., some images contain dense objects concentrated in specific areas, while some only have a few objects scattered throughout the images.
Stemming on the above-mentioned weakness, we proposed a novel DETR-like method named DQ-DETR which mainly focuses on dynamically adapting the numbers of queries and enhancing the position of queries to locate the tiny objects precisely. In this work, we propose a dynamic query selection module for adaptively choosing different numbers of object queries in DETR’s decoder stage, resulting in fewer FP in sparse images and fewer FN in dense images. Moreover, we generate the density maps and estimate the number of instances in an image by the categorical counting module. The number of object queries is adjusted based on the predicted counting number. In addition, we aggregate the density maps with the visual feature from the transformer encoder to reinforce the foreground features, enhancing the spatial information for tiny objects. The strengthened visual feature will be further used to improve the positional information of object queries. As such, we can simultaneously handle the images with few and crowded tiny objects by dynamically adjusting the number and position of object queries used in the decoder.
@@ 2 Related work
2.1 DETR-like Methods.
DETR [3] proposes an end-to-end object detection framework based on the transformer named DETR (DEtection TRansformer), where the transformer encoder extracts instance-level features from an image, and the transformer decoder uses a set of learnable queries to probe and pool features from images. While DETR achieves comparable results with the previous classical CNN-based detectors [1, 2], it suffers severely from the problem of slow training convergence, needing 500 epochs of training to perform well. Many follow-up works have attempted to address the slow training convergence of DETR from different perspectives.
@@
2.2 Tiny Object Detection.
@@
However, these methods highly rely on the predefined threshold, which is unstable to a different dataset. Instead, our DQ-DETR uses a counting-guided feature enhancement module to improve the visual features with more spatial information about tiny objects’ scale and location. Also, although the above-mentioned DETR-like methods have improved the formulation of queries, they are not specially designed to detect tiny objects. For example, the object queries in [7, 13, 3, 5] are learned from the training data and remain the same for different input images. Our proposed DQ-DETR is the first DETR-like model that focuses on tiny object detection. DQ-DETR dynamically adjusts the number of object queries and enhances the position information of queries in order to detect tiny objects precisely under imbalanced situations in aerial datasets.
@@ 3.3 Categorical Counting Module
3.3.2 Counting Number Classification.
Lastly, we estimate the counting number $N$, i.e., the number of instances per image, by a classification head and categorize them into four levels, which are $N\leq 10$, $10<N\leq 100$, $100<N\leq 500$, and $N>500$. The classification head consists of two linear layers. Further, the numbers 10, 100, and 500 are selected based on the dataset’s characteristics, i.e., the mean and standard deviation of the number of instances $N$ per image. Notably, we do not use the regression head as in the traditional crowd-counting methods, which regresses the counting number to a specific number. We attribute the reason to the drastic difference of the number of instances in each image, where $N$ ranges from 1 to 2267 in different images of AI-TOD-V2. It is difficult to regress an accurate number, hurting the detection performance.
@@ 3.4 Counting-Guided Feature Enhancement Module (CGFE)
The EMSV feature will be refined with the density map from the categorical counting module through the proposed Counting-Guided Feature Enhancement Module (CGFE) to improve the spatial information of tiny objects. The refined features will further be used to enhance the position information of queries. This module comprises a spatial cross-attention and a channel attention operations.
3.4.1 Spatial cross-attention map.
To utilize the abundant spatial information in the density map ${F}_{c}$, a 2-D cross-spatial attention is calculated. We employ a $1\times 1$ convolution layers to down-sample the density map ${F}_{c}$, creating multi-scale counting feature maps ${F}_{c,i}\in\{1,2,\dots,l\}$ to in line with the shape of each layer of encoder’s multi-scale feature maps $S_{i}\in\{1,2,\dots,l\}$. Subsequently, we first apply average pooling (AvgP.) and max pooling (MaxP.) on each layer of multi-scale counting features ${F}_{c,i}\in\mathbb{R}^{b\times 256\times h\times w}$ along the channel axis. Then, the two pooling features $\mathbb{R}^{b\times 1\times h\times w}$ are concatenated and sent into a 7x7 convolution layer followed by a Sigmoid function to produce spatial attention map ${W}_{s}\in\mathbb{R}^{b\times 1\times h\times w}$. We formulate this process in Eq. 1.
@@
3.4.2 Channel attention map.
After the spatial cross-attention, we further apply 1-D channel attention to the spatial-intensified features ${E}_{i}$, exploiting the inter-channel relationship of features. Specifically, we first apply average pooling and max pooling on each layer of ${E}_{i}\in\mathbb{R}^{b\times 256\times h\times w}$ along the spatial dimension. Next, the two pooling features $\mathbb{R}^{b\times 256\times 1\times 1}$ are sent into a shared MLP and merged together with element-wise addition to create channel attention map ${W}_{c,i}$. Finally, the channel attention map ${W}_{c,i}\in\mathbb{R}^{b\times 256\times 1\times 1}$ is multiplied with original $E_{i}\in\mathbb{R}^{b\times 256\times h\times w}$ and further get the counting-guided intensified feature maps ${F}_{t}$. The formulas are defined in Eq. 3 and Eq. 4:
${W}_{c,i}=\sigma(MLP(AvgP.({E}_{i}))+MLP(MaxP.({E}_{i}))),$ (3)
@@ 4 Experiments
4.1 Dataset
AI-TOD-V2. This dataset includes 28,036 aerial images with 752,745 annotated object instances. There are 11,214 images for the train set, 2,804 for the validation set, and 14,018 for the test set. The average object size in AI-TOD-V2 is only 12.7 pixels, with 86% of objects in the dataset smaller than 16 pixels, and even the largest object is no bigger than 64 pixels. Also, the number of objects in an image can vary enormously from 1 to 2667, where the average number of objects per image is 24.64 with a standard deviation of 63.94.
Evaluation Metric. We use the AP (Average Precision) metric with a max detection number of 1500 to evaluate the performance of our proposed method. Specifically, AP means the average value from $\mathrm{AP_{0.5}}$ to $\mathrm{AP_{0.95}}$ , with IoU interval of 0.05. Moreover, $\mathrm{AP_{vt}}$, $\mathrm{AP_{t}}$, $\mathrm{AP_{s}}$, and $\mathrm{AP_{m}}$ are for very tiny, tiny, small, and medium scale evaluation in AI-TOD [21].
4.2 Implementation Details
Based on the DETR-like structure, we use a 6-layer transformer encoder, a 6-layer transformer decoder with 256 as the hidden dimension, and a ResNet50 as our CNN backbone. Furthermore, we train our model for 24 epochs with Adam optimizer and weight decay 0.0001 using 2 NVIDIA 3090 GPUs. The batch size is set to 1 due to memory constraints. The same random crop and scale augmentation strategies are applied following DETR [3]. Also, we apply a two-stage training scheme. We first train the categorical counting module to achieve more stable results for the number of queries in the transformer decoder. After stabilizing the counting result, we add the counting-guided feature enhancement module into training to refine the encoder’s visual features with density maps.
Method #Epochs Backbone $\mathrm{AP}$ $\mathrm{AP_{50}}$ $\mathrm{AP_{75}}$ $\mathrm{AP_{vt}}$ $\mathrm{AP_{t}}$ $\mathrm{AP_{s}}$ $\mathrm{AP_{m}}$
@@ 4.2 Implementation Details
4.3 Main Results on AI-TOD-V2
Table 2 shows our main results on the AI-TOD-V2 test split. We compare the performances of our DQ-DETR with strong baselines, including both CNN-based and DETR-like methods. All CNN-based methods except YOLOv3 use ResNet50 with feature pyramid network (FPN) [25]. Moreover, since there is no previous research of DETR-like models on tiny object detection, our DQ-DETR is the first DETR-like model that focuses on detecting tiny objects. We re-implement a series of DETR-like models on AI-TOD-V2, and all DETR-like methods except DETR use 5 scales of feature maps with deformable attention [4]. For 5-scale feature maps, features are extracted from stages 1, 2, 3, and 4 of the backbone, and add the extra feature by down-sampling the output of stage 4.
The results are summarized in Table 2, our proposed DQ-DETR achieves the best result 30.2 AP compared with other state-of-the-art methods, including CNN-based and DETR-like methods. Also, DQ-DETR surpasses the baseline by 20.5%, 20.6%, 14.1%, and 12.3% in terms of $\mathrm{AP_{vt}}$, $\mathrm{AP_{t}}$, $\mathrm{AP_{s}}$, $\mathrm{AP_{m}}$. The performance gain is greater on $\mathrm{AP_{vt}}$, and $\mathrm{AP_{t}}$, and our DQ-DETR outperforms the advanced series of DETR-like models on AI-TOD-V2. We credit the performance gains for the following reasons: (1) DQ-DETR fuses the transformer visual features with a density map from the categorical counting module to improve the positional information of object queries, which makes the queries more suitable for localizing tiny objects. (2) Our dynamic query selection adaptively chooses an adequate number of object queries used for the detection task and can handle the images with either few or crowded objects.
Model $\mathrm{AP}$
$\mathrm{AP_{50}}$
$\mathrm{AP_{75}}$
Faster R-CNN [1] 21.4
40.7
19.9
Cascade R-CNN [23] 22.6
38.8
23.2
Yolov5 [26] 24.1
44.1
22.3
CEASC [27] 28.7
50.7
24.7
SDP [28] 30.2
52.5
28.4
DINO-DETR* [6] 35.8
58.3
36.8
DQ-DETR (Ours) 37.0 60.9 37.9
4.4 Results on VisDrone
Besides AI-TOD-V2, we also conduct experiments on the VisDrone [29] dataset to demonstrate the effectiveness of our model DQ-DETR. VisDrone. This dataset includes 14,018 drone-shot images, with 6,471 images for the train set, 548 for the validation set, and 3,190 for the test set. There are 10 categories, and the image resolution is 2000 $\times$ 1500 pixels. Also, the images are diverse in a wide range of aspects, including objects (pedestrians, vehicles, bicycles, etc.) and density (sparse and crowded scenes), where the average number of objects per image is 40.7 with a standard deviation of 46.41.
4.5 Ablation Study
Categorical counting, counting-guided feature enhancement, and dynamic query selection are the newly proposed contributions. We conduct a series of ablation studies to verify the effectiveness of each component proposed in this paper. DINO-DETR is chosen as the comparing DETR-like baseline.
4.5.1 Main ablation experiment.
Table 4 shows the performance of our contributions separately on AI-TODv2. The results demonstrate that each component in DQ-DETR contributes to performance improvement. We attain an improved +2.2 AP over the baseline with the categorical counting module and dynamic query selection. Furthermore, with feature enhancement refining the encoder’s feature, it gains an extra improvement of +4.3, +2.6, +5.2 on $\mathrm{AP}$, $\mathrm{AP_{vt}}$, and $\mathrm{AP_{t}}$ over the baseline. Besides, the experiment with the counting module and feature enhancement together but without dynamic query selection further shows that introducing an additional counting-guided feature-enhancing task improves performance, even when query numbers remain static. Consequently, we prove the power of each component in DQ-DETR on AI-TOD-V2.
Categorical Counting (CC) Dynamic Query Selection (DQS) Feature Enhancement (FE) $\mathrm{AP}$ $\mathrm{AP_{vt}}$ $\mathrm{AP_{t}}$ $\mathrm{AP_{s}}$ $\mathrm{AP_{m}}$
@@ 4.5 Ablation Study
28.1 12.3 27.8 34.6 44.1
29.1 14.4 29.3 35.2 44.1
30.2 15.3 30.5 36.5 44.6
4.5.2 Ablation of DQ-DETR with different number of instances in images.
We explore our DQ-DETR’s performance under different numbers of instances in an image. We classify the AI-TOD-V2 dataset into 4 levels based on the number of instances $N$ in the image as in the categorical counting module, i.e., $N\leq 10$, $10<N\leq 100$, $100<N\leq 500$, and $500<N$. Our proposed DQ-DETR’s performance is analyzed under these four situations. The results are shown in Table 5 compared with DINO-DETR as the baseline. Our DQ-DETR dynamically adjusts the number of object queries based on the number of instances in the image. while DINO-DETR always uses 900 queries in all situations.
We can observe that in the situations of $N\leq 10$, and $10<N\leq 100$, our DQ-DETR uses fewer numbers of queries and outperforms the baseline by 16%, and 16.4% in terms of $\mathrm{AP}$. The performances in terms of $\mathrm{AP_{vt}}$, $\mathrm{AP_{t}}$ surpass the baseline by 19.8%, and 20.8% as well. Moreover, it is noteworthy that DINO-DETR performs poorly when $N>500$. Under these circumstances, there might be over 900 instances in some of the images, which is beyond the detection capability of DINO-DETR. In dense images, the detection limitation of DINO-DETR with only 900 queries will lead to many objects undetected (FN), resulting in a lower $\mathrm{AP}$. Our DQ-DETR dynamically selects more queries for dense images, remarkably surpassing the baseline by 42.1% in terms of $\mathrm{AP_{vt}}$.
Model #objects in image #Query $\mathrm{AP}$ $\mathrm{AP_{50}}$ $\mathrm{AP_{75}}$ $\mathrm{AP_{vt}}$ $\mathrm{AP_{t}}$ $\mathrm{AP_{s}}$ $\mathrm{AP_{m}}$
@@
4.5.3 Ablation of Categorical Counting Module.
#Objects in image Accuracy(%) AP (DQ-DETR) AP (Baseline) #Sample
@@
Method $\mathrm{AP}$ $\mathrm{AP_{vt}}$ $\mathrm{AP_{t}}$ $\mathrm{AP_{s}}$ $\mathrm{AP_{m}}$
@@
Table 7 compares our DQ-DETR’s performance of using classification or regression in the categorical counting module. The traditional crowd-counting methods usually regress the predicted counting number to a specific value. However, in our study, we use a classification head instead. This experiment demonstrates our DQ-DETR performance with these two methods. For the classification task, we classify the images into 4 classes and apply different numbers of queries in the transformer decoder, as we mentioned in the previous section. For the regression task, we regress an integer directly to predict the number of objects in the images and select the object queries corresponding to the predicted result.
The results demonstrate that using regression as a counting method performs extremely poorly. We impute the drastic performance drop for the following reasons: (1) It is challenging to regress an accurate number since the number of instances per image may vary significantly from 1 to 2267 in the AI-TOD-V2 dataset. (2) The unstable regression result will largely affect the number of queries used in the transformer decoder and cause the DETR model hard to converge. Owing to the above reasons, we believe that classifying how many objects exist in the image into different levels is a simpler way in contrast to regression. Thus, we take classification instead of regression as a preferable method in our proposed categorical counting module.